# PerlDiff: Controllable Street View Synthesis Using Perspective-Layout Diffusion Models

### v1

@@ Abstract
Controllable generation is considered a potentially vital approach to address the challenge of annotating 3D data, and the precision of such controllable generation becomes particularly imperative in the context of data production for autonomous driving. Existing methods focus on the integration of diverse generative information into controlling inputs, utilizing frameworks such as GLIGEN or ControlNet, to produce commendable outcomes in controllable generation. However, such approaches intrinsically restrict generation performance to the learning capacities of predefined network architectures. In this paper, we explore the integration of controlling information and introduce PerlDiff (Perspective-Layout Diffusion Models), a method for effective street view image generation that fully leverages perspective 3D geometric information. Our PerlDiff employs 3D geometric priors to guide the generation of street view images with precise object-level control within the network learning process, resulting in a more robust and controllable output. Moreover, it demonstrates superior controllability compared to alternative layout control methods. Empirical results justify that our PerlDiff markedly enhances the precision of generation on the NuScenes and KITTI datasets. Our codes and models are publicly available at https://github.com/LabShuHangGU/PerlDiff.
@@ 1 Introduction
To mitigate the issue of data scarcity, the adoption of generative technologies [32, 5, 9] has proven practical for reversing the order of data annotation. The paradigm of this approach is to use the collected annotation as controlling information to generate the corresponding lifelike images depicting urban street scenes. By implementing this strategy, it is possible to dramatically lower the costs associated with data annotation while also facilitating the generation of extensive long-tail datasets, subsequently leading to improvements in the perception model’s performance. Pioneering research, exemplified by BEVGen [30], harnesses the capabilities of autoregressive transformers [32, 5] to render detailed visualizations of street scenes. Additionally, subsequent studies such as BEVControl [38] and MagicDrive [6] employ diffusion-based techniques, including GLIGEN [13] and ControlNet [39], to integrate controlling information through a basic cross-attention mechanism. However, these methods simply extract integrated conditional features from controlling information to adjust the generation process and are limited in making full use of detailed geometric layout information for accurate attention map manipulation. While these techniques make strides towards meeting the requirements for generating 3D annotations, Fig. 1 demonstrates that there remains significant potential for improvement in object controllability.
Figure 1: PerlDiff enhances controllability over BEVControl* using geometric priors. First row: Adjusting BEV 3D annotation yaw changes the vehicle angles in generated images. Second row: Shifting annotated BEV coordinates moves vehicles forward in the scene. Third row: Morphing the object into changing vehicle types such as a car, a truck, and a bus increases the object size.
@@
To this end, in this paper, we introduce the perspective-layout diffusion models (PerlDiff), a novel method specifically designed to enable precise control over street view image generation at the object level. In addition to extracting integrated conditional features from controlling condition information, i.e. 3D annotations, our PerlDiff model explicitly renders perspective layout masking maps as geometric priors. Subsequently, a Perl-based controlling module (Perl-CM) is proposed to leverage the geometric priors, i.e. perspective layout masking maps. Within Perl-CM, an innovative Perl-based cross-attention mechanism is utilized to accurately guide the generation of each object with their corresponding condition information. We integrate Perl-CM into the pre-trained Stable Diffusion model [24] and fine-tune it on our training dataset. Consequently, our PerlDiff incorporates the formidable generative capabilities of Stable Diffusion with the finely detailed geometric priors of perspective layouts, effectively harnessing their respective strengths for precise object-level image synthesis. Overall, our PerlDiff is capable of generating precise, controllable street view images while also maintaining high fidelity (see Section 3 for details).
@@
The main contributions of this paper are summarized as following three-folds: (i) We present PerlDiff, a newly developed framework crafted to synthesize street view images from user-defined 3D annotations. Our PerlDiff carefully orchestrates the image generation process at the object level by leveraging perspective layout masks as geometric priors. (ii) We propose a Perl-based cross-attention mechanism that utilizes perspective layout masking maps from 3D annotations to enhance the underlying cross-attention mechanism within Perl-CM. This method enables precise control over the street view image generation process by integrating road geometry and object-specific information derived from BEV annotations. (iii) Our PerlDiff method attains state-of-the-art performance on the NuScenes [1] and KITTI [7] dataset compared to existing methods, markedly enhancing detection and segmentation outcomes for synthetic street view images. Furthermore, it holds the potential to function as a robust traffic simulator in the future.
@@ 2 Related Work
Geometric Constraints in Image Generation. Incorporating geometric priors into image synthesis has been explored to a lesser extent. Work on 3D-aware image generation [20, 21] suggests the feasibility of integrating geometric information into generative processes to improve spatial coherence. Nevertheless, these methods often rely on complex 3D representations and may not be directly applied to diffusion model frameworks. Recently, BoxDiff [37] reveals a spatial correspondence between the attention map produced by the diffusion model and the corresponding generated image. During the testing phase, the geometric configuration of the attention map’s response values is adjusted to yield a more precise image generation. ZestGuide [3] introduces a loss function that enforces a geometric projection onto the attention map, further refining the shape of the attention map’s response values to closely approximate the geometric projection of the control information during the inference stage. However, utilizing text prompts to facilitate the generation of complex urban environment layouts poses inherent challenges, owing to the need for crafting intricate prompts to accurately depict urban environments. Furthermore, modifying the cross-attention map to impose strict constraints during the denoising phase of inference can disrupt the intrinsic relationships, leading to a suboptimal approach to synthesizing controllable images. In contrast, our PerlDiff incorporates geometric prerequisites as training priors to guide the generation of street view images, offering a more effective solution.
@@ 3 Controllable Street View Generation Based on Perspective Layout
In this paper, we introduce PerlDiff, depicted in Fig. 2, which is designed to enable controllable multi-view street scene generation using BEV annotations. Specifically, PerlDiff leverages perspective projection information from BEV annotations as controlling condition within the training regimen and utilizes perspective layout masks as geometric priors, enabling accurate guidance in object generation. In the following sections, we delineate the process of encoding the controlling condition information from the BEV annotations in Section 3.1. Additionally, we explain how incorporating perspective layout knowledge ensures object controllability in street view image generation in Section 3.2.
Figure 2: Overview of PerlDiff framework for multi-view street image generation. our PerlDiff utilizes perspective layout masking maps derived from BEV annotations to integrate scene information and object bounding boxes. Perl-CM is responsible for integrating control information through employing Perl-based cross-attention (Scene & Object) mechanism, using Perl masking map (road & box) as geometric priors to guide object-level image generation with high precision. View cross-attention ensures consistency across multiple views, while Text cross-attention integrates textual scene descriptions to facilitate further adjustments.
@@ 3.1 Controlling Conditions Encoding
Given BEV annotations of a street scene, our goal is to generate multi-view street images. To be more specific, for controllable street view image generation, we extract not only scene information (i.e., textual scene descriptions $\mathbf{S}_{d}$ and street maps $\mathbf{M}$ revealing features such as road markings and obstacles) but also object information (i.e., bounding box parameters $\mathbf{P}$ and the associated object category $\mathbf{Y}$) from BEV annotations as controlling conditions. These controlling conditions encompass rich semantic and geometric information, so establishing a robust encoding method to utilize this information is essential for generating street view images. Hereafter, we present our controlling condition encoding approach. We omit the details of multi-view perspectives for notational simplicity.
@@
Perl Object Information encapsulates perspective geometric data alongside object category information, which stems from the projections of annotated BEV 3D boxes. This element plays a pivotal role in enabling inverse 3D labeling. Through projecting BEV 3D annotations onto their corresponding perspective images, we ascertain eight 2D corner points for each bounding box within a single image, denoted as $\mathbf{P}_{g}\in\mathbb{R}^{M\times 2\times 8}$, where $M$ represents the maximum number of bounding boxes and eight corresponds to the number of corners per bounding box. In conjunction with the object’s categorical text $\mathbf{P}_{c}=\{\mathbf{P}_{c_{i}}\}_{i=1}^{M}$, we derive the encoded box geometric features $\mathbf{H}_{g}\in\mathbb{R}^{M\times C}$ and the box categorical features $\mathbf{H}_{c}\in\mathbb{R}^{M\times C}$, which are illustrated as follows:
@@
where $\mathcal{F}(\cdot)$ is the Fourier embedding [18] function, $\varphi(\cdot)$ represents the pre-trained text embedding encoder of CLIP [23] and $C$ representing the dimension of features. Furthermore, we concatenate the encoded geometric features $\mathbf{H}_{g}$ and categorical features $\mathbf{H}_{c}$ and subsequently pass the concatenated vector through a Multilayer Perceptron (MLP)  [31] $\mathbf{F}$, to achieve feature fusion $\mathbf{H}_{b}\in\mathbb{R}^{M\times C}$. The resulting fused box feature representation is given by:
@@
We subsequently input the encoded conditions into the denoising diffusion model to guide the generation process. This is achieved utilizing Perl-based cross-attention mechanism that incorporates Perl masking maps, as detailed below.
@@ 3.2 Object Controllability via Perl-based Controlling Module
Perl-based Controlling Module (Perl-CM) is responsible for integrating controlling condition information, which encompasses both the Perl scene and object information, into the latent feature maps of noisy street view images. This integration is primarily achieved via the scene and object Perl-based cross-attention mechanism. Initially, this mechanism assigns initial values to the attention maps, under the guidance of road and box Perl masking maps. Throughout the training of the network, these values are optimized to ensure that the response of the attention map accurately corresponds to the regions where the objects are located. Subsequently, the information from both the road map and the bounding box are sequentially integrated into the noise street view image. To more effectively integrate Perl scene and object information, a gating operation is used, similar to the method employed in GLIGEN  [13], which dynamically adjusts the contribution of condition information according to the adaptive process. To ensure multi-view consistency, View Cross-attention leverages information from the immediate left and right views for uniformity across various perspectives. Additionally, Text Cross-attention manipulates the weather and lighting conditions of street scenes using textual scene description. The details of the Perl-CM process are delineated in the Algo. 1.
@@
Figure 3: Visualization of cross-attention maps reveals perceptual congruency with the generated image. BEVControl* produces disorganized and vague attention maps, which result in inferior image quality. Conversely, our PerlDiff method fine-tunes the response within the attention maps, resulting in more accurate control information at the object level and improved image quality. See more qualitative examples in the Appendix D.
@@
$\mathcal{A}_{s}=\textit{softmax}(\lambda_{s}\cdot\mathcal{M}_{s}+\mathbf{Z}\mathbf{H}_{m}^{T}/\sqrt{d}),\ \ \ \ \mathcal{A}_{b}=\textit{softmax}(\lambda_{b}\cdot\mathcal{M}_{b}+\mathbf{Z}\mathbf{H}_{b}^{T}/\sqrt{d})$ (5)
@@
where $\lambda_{s}$ and $\lambda_{b}$ are weight parameters that control the influence of the masking map, $d$ denotes the dimensionality, and $\mathbf{Z}\in\mathbb{R}^{HW\times C}$ is the noisy street view image. Here, $\mathcal{A}_{s}\in\mathbb{R}^{HW\times 1}$ characterizes the association between the road map and the noisy image, while $\mathcal{A}_{b}\in\mathbb{R}^{HW\times M}$ clarifies the relationship between the conditions of the object bounding box and the noisy image. The final noisy street view image is synthesized through an attention-based aggregation mechanism enhanced by a residual connection, which can be expressed as:
$\mathbf{Z}=\gamma_{s}\cdot\mathcal{A}_{s}\mathbf{H}_{m}+\mathbf{Z},\qquad\qquad\ \ \ \ \mathbf{Z}=\gamma_{b}\cdot\mathcal{A}_{b}\mathbf{H}_{b}+\mathbf{Z},$ (6)
@@
where $\gamma_{s}$ and $\gamma_{b}$ represent learnable parameters modulating the influence of the respective conditions.
@@
View Cross-attention is corroborated by preliminary works such as BEVControl [38], MagicDrive [6], DrivingDiffusion [12], and Panacea [36], plays an instrumental role in facilitating the synthesis of images that maintain visual consistency across varying camera perspectives. For additional information, please see the Appendix C.3.
@@
Text Cross-attention enhances the Stable Diffusion [24] model’s ability to modulate street scenes through textual scene description. This capability is crucial for dynamically adapting the rendering of street scenes to accommodate various lighting and weather conditions. By integrating detailed textual scene description, our PerlDiff can effectively alter visual elements such as illumination and atmospheric effects, ensuring that the generated images reflect the specified conditions accurately. Please refer to Fig. 4 for qualitative examples of this enhancement.
Algorithm 1 Perl-based Controlling Module (Perl-CM)
2:Output: Updated $\mathbf{Z}$
5:$\mathbf{Z}\leftarrow\gamma_{s}\cdot\mathcal{A}_{s}\mathbf{H}_{m}+\mathbf{Z}$
6:$\mathcal{A}_{b}\leftarrow\textit{softmax}(\lambda_{b}\cdot\mathcal{M}_{b}+\mathbf{Z}\mathbf{H}_{b}^{T}/\sqrt{d})$
8:$\mathbf{Z}\leftarrow\gamma_{b}\cdot\mathcal{A}_{b}\mathbf{H}_{b}+\mathbf{Z}$
9:$\mathbf{Z}\leftarrow\mathbf{Z}_{in}+\mathcal{C}(\mathbf{Z}_{in},\mathbf{Z}_{l},\mathbf{Z}_{l})+\mathcal{C}(\mathbf{Z}_{in},\mathbf{Z}_{r},\mathbf{Z}_{r})$
10:// maintain visual consistency via view cross-attention

### v4

@@ Abstract
Controllable generation is considered a potentially vital approach to address the challenge of annotating 3D data, and the precision of such controllable generation becomes particularly imperative in the context of data production for autonomous driving. Existing methods focus on the integration of diverse generative information into controlling inputs, utilizing frameworks such as GLIGEN or ControlNet, to produce commendable outcomes in controllable generation. However, such approaches intrinsically restrict generation performance to the learning capacities of predefined network architectures. In this paper, we explore the innovative integration of controlling information and introduce PerLDiff (Perspective-Layout Diffusion Models), a novel method for effective street view image generation that fully leverages perspective 3D geometric information. Our PerLDiff employs 3D geometric priors to guide the generation of street view images with precise object-level control within the network learning process, resulting in a more robust and controllable output. Moreover, it demonstrates superior controllability compared to alternative layout control methods. Empirical results justify that our PerLDiff markedly enhances the precision of controllable generation on the NuScenes and KITTI datasets.
@@ 1 Introduction
Figure 1: PerLDiff enhances controllability over BEVControl* and MagicDrive using geometric priors. Top: Demonstrates object controllability by adjusting the 3D annotation yaw by 90 degrees. Bottom: Shows scene controllability through the alignment of the street map with the generated image. Regions highlighted by rectangles indicate areas where the generated images fail to achieve control and alignment with ground truth conditions.
@@
To mitigate issue of data scarcity, the adoption of generative technologies [32, 5, 10] has proven practical for reversing the order of data annotation. The paradigm of this approach is to use the collected annotation as controlling information to generate the corresponding lifelike images depicting urban street scenes. By implementing this strategy, it is possible to dramatically lower the costs associated with data annotation while also facilitating the generation of extensive long-tail datasets, subsequently leading to improvements in the perception model’s performance. Pioneering research, exemplified by BEVGen [30], harnesses the capabilities of autoregressive transformers [32, 5] to render detailed visualizations of street scenes. Additionally, subsequent studies such as BEVControl [38] and MagicDrive [6] employ diffusion-based techniques, including GLIGEN [13] and ControlNet [39], to integrate controlling information through a basic cross-attention mechanism. However, these methods simply extract integrated conditional features from controlling information to adjust the generation process and are limited in making full use of detailed geometric layout information for accurate attention map manipulation. While these techniques make strides towards meeting the requirements toward generating data from 3D annotations, Fig. 1 demonstrates that there remains significant potential for improvement in scene and object controllability.
@@
To this end, in this paper, we introduce the perspective-layout diffusion models (PerLDiff), a novel method specifically designed to enable precise control over street view image generation at the object level. In addition to extracting integrated conditional features from controlling condition information, i.e. 3D annotations, our PerLDiff explicitly renders perspective layout masking maps as geometric priors. Subsequently, a PerL-based controlling module (PerL-CM) is proposed to leverage the geometric priors, i.e. perspective layout masking maps. Within PerL-CM, an innovative PerL-based cross-attention mechanism is utilized to accurately guide the generation of each object with their corresponding condition information. We integrate PerL-CM into the pre-trained Stable Diffusion model [24] and fine-tune it on our training dataset. Consequently, our PerLDiff incorporates the formidable generative capabilities of Stable Diffusion with the finely detailed geometric priors of perspective layouts, effectively harnessing their respective strengths for precise object-level image synthesis. Overall, our PerLDiff is capable of generating precise, controllable street view images while also maintaining high fidelity (see Section 3 for details).
The main contributions of this paper are summarized as following three-folds:
  • We present PerLDiff, a newly developed framework crafted to synthesize street view images from user-defined 3D annotations. Our PerLDiff carefully orchestrates the image generation process at the object level by leveraging perspective layout masks as geometric priors.
@@
  • We propose a PerL-based cross-attention mechanism that utilizes perspective layout masking maps from 3D annotations to enhance the underlying cross-attention mechanism within PerL-CM. This method enables precise control over the street view image generation process by integrating road geometry and object-specific information derived from 3D annotations.
  • Our PerLDiff method attains state-of-the-art performance on the NuScenes [1] and KITTI [7] dataset compared to existing methods, markedly enhancing detection and segmentation outcomes for synthetic street view images. Furthermore, it holds the potential to function as a robust traffic simulator in the future.
@@ 2 Related Work
Figure 2: Overview of PerLDiff framework for multi-view street image generation (multi-view dimension omitted for simplicity). PerLDiff utilizes perspective layout masking maps derived from 3D annotations to integrate scene information and object bounding boxes. PerL-CM is responsible for integrating control information through employing PerL-based cross-attention (Scene & Object) mechanism, using PerL masking map (road & box) as geometric priors to guide object-level image generation with high precision. View cross-attention ensures consistency across multiple views, while Text cross-attention integrates textual scene descriptions to facilitate further adjustments.
@@
Geometric Constraints in Image Generation. Incorporating geometric priors into image synthesis has been explored to a lesser extent. Work on 3D-aware image generation [20, 21] suggests the feasibility of integrating geometric information into generative processes to improve spatial coherence. Nevertheless, these methods often rely on complex 3D representations and may not be directly applied to diffusion model frameworks. Recently, BoxDiff [37] reveals a spatial correspondence between the attention map produced by the diffusion model and the corresponding generated image. During the testing phase, the geometric configuration of the attention map’s response values is adjusted to yield a more precise image generation. ZestGuide [3] introduces a loss function that enforces a geometric projection onto the attention map, further refining the shape of the attention map’s response values to closely approximate the geometric projection of the control information during the inference stage. However, utilizing text prompts to facilitate the generation of complex urban environment layouts poses inherent challenges, owing to the need for crafting intricate prompts to accurately depict urban environments. Furthermore, modifying the cross-attention map to impose strict constraints during the denoising phase of inference can disrupt the intrinsic relationships, leading to a suboptimal approach to synthesizing controllable images. In contrast, PerLDiff incorporates geometric prerequisites as training priors to guide the generation of street view images, offering a more effective solution.
@@ 3 Controllable Street View Generation Based on Perspective Layout
In this paper, we introduce PerLDiff, depicted in the Fig. 2, which is designed to enable controllable multi-view street scene image generation using 3D annotations. Specifically, our PerLDiff leverages perspective projection information from 3D annotations as controlling condition within the training regimen and utilizes perspective layout masks as geometric priors, enabling accurate guidance in object generation. In the following sections, we delineate the process of encoding the controlling condition information from the 3D annotations in the Section 3.1. Additionally, we explain how incorporating the perspective layout knowledge to ensure scene and object controllability in street view image generation in the Section 3.2.
@@ 3.1 Controlling Conditions Encoding
Given 3D annotations of a street scene, our goal is to generate multi-view street images. To be more specific, for controllable street view image generation, we extract not only scene information (i.e., textual scene descriptions $\mathbf{S}_{d}$ and street maps $\mathbf{M}$ revealing features such as road markings and obstacles) but also object information (i.e., bounding box parameters $\mathbf{P}$ and the associated object category $\mathbf{Y}$) from 3D annotations as controlling conditions. These controlling conditions encompass rich semantic and geometric information, so establishing a robust encoding method to utilize this information is essential for generating street view images. Hereafter, we present our controlling condition encoding approach. For simplicity, we omit the details of multi-view perspectives.
@@
PerL Object Information encapsulates perspective geometric data alongside object category information, which stems from the projections of annotated 3D boxes. Through projecting 3D annotations onto their corresponding perspective images, we ascertain eight 2D corner points for each bounding box within a single image, denoted as $\mathbf{P}_{g}\in\mathbb{R}^{M\times 2\times 8}$, where $M$ represents the maximum number of bounding boxes and eight corresponds to the number of corners per bounding box. In conjunction with the object’s categorical text $\mathbf{P}_{c}=\{\mathbf{P}_{c_{i}}\}_{i=1}^{M}$, we derive the encoded box geometric features $\mathbf{H}_{g}\in\mathbb{R}^{M\times C}$ and the box categorical features $\mathbf{H}_{c}\in\mathbb{R}^{M\times C}$, which are illustrated as follows:
@@
where $\mathcal{F}(\cdot)$ is the Fourier embedding function [18], $\varphi(\cdot)$ represents the pre-trained text embedding encoder of CLIP [23] and $C$ representing the dimension of features. Furthermore, we concatenate the encoded geometric features $\mathbf{H}_{g}$ and categorical features $\mathbf{H}_{c}$ and subsequently pass the concatenated vector through a Multilayer Perceptron (MLP) [31] $\mathbf{F}$, to achieve feature fusion $\mathbf{H}_{b}\in\mathbb{R}^{M\times C}$. The resulting fused box feature representation is given by:
@@
We subsequently input the encoded conditions into the denoising diffusion model to guide the generation process. This is achieved utilizing PerL-based cross-attention that incorporates PerL masking maps, as detailed below.
Figure 3: Visualization of cross-attention maps reveals perceptual congruency with the generated image. BEVControl* produces disorganized and vague attention maps, which result in inferior image quality. Conversely, our PerLDiff method fine-tunes the response within the attention maps, resulting in more accurate control information at the object level and improved image quality. Please see more qualitative examples in the Appendix.
@@ 3.2 Object Controllability via PerL-based Controlling Module
PerL-based Controlling Module (PerL-CM) is responsible for integrating controlling condition information, which encompasses both the PerL scene and object information, into the latent feature maps of noisy street view images. This integration is primarily achieved via the scene and object PerL-based cross-attention mechanism. Initially, this mechanism assigns initial values to the attention maps, under the guidance of road and box PerL masking maps. Throughout the training of the network, these values are optimized to ensure that the response of the attention map accurately corresponds to the regions where the objects are located. Subsequently, the information from both the road map and the bounding box are sequentially integrated into the noise street view image. To more effectively integrate PerL scene and object information, a gating operation similar to GLIGEN [13] is used, which dynamically adjusts the contribution of condition information according to the adaptive process. To ensure multi-view consistency, View Cross-attention leverages information from the immediate left and right views for uniformity across various perspectives. Additionally, Text Cross-attention manipulates the weather and lighting conditions of street scenes using textual scene description. The details of the PerL-CM process are delineated in the Appendix.
@@
$\mathcal{A}_{s}=\textit{softmax}(\lambda_{s}\cdot\mathcal{M}_{s}+\mathbf{Z}\mathbf{H}_{m}^{T}/\sqrt{d}),$ (5)
$\mathcal{A}_{b}=\textit{softmax}(\lambda_{b}\cdot\mathcal{M}_{b}+\mathbf{Z_{s}}\mathbf{H}_{b}^{T}/\sqrt{d})$ (6)
@@
where $\lambda_{s}$ and $\lambda_{b}$ are weight parameters that control the influence of the masking map, $d$ denotes the dimensionality, and $\mathbf{Z}$, $\mathbf{Z}_{b}$, $\mathbf{Z}_{s}\in\mathbb{R}^{HW\times C}$ represent different noisy street view images. Here, $\mathcal{A}_{s}\in\mathbb{R}^{HW\times 1}$ characterizes the association between the road map and the noisy image, while $\mathcal{A}_{b}\in\mathbb{R}^{HW\times M}$ clarifies the relationship between the conditions of the object bounding box and the noisy image. For visualization purposes, as shown in Fig. 3, we average $\mathcal{A}_{b}$ along the second dimension and merge all object attention maps into one representation at step 50 of the DDIM process in the last block of U-Net. The final noisy street view image is synthesized through an attention-based aggregation mechanism enhanced by a residual connection, which can be expressed as:
$\mathbf{Z_{s}}=\gamma_{s}\cdot\mathcal{A}_{s}\mathbf{H}_{m}+\mathbf{Z},\qquad\mathbf{Z_{b}}=\gamma_{b}\cdot\mathcal{A}_{b}\mathbf{H}_{b}+\mathbf{Z_{s}},$ (7)
@@
where $\gamma_{s}$ and $\gamma_{b}$ represent learnable parameters modulating the influence of respective conditions.
@@
View Cross-attention is corroborated by preliminary works such as BEVControl [38], MagicDrive [6], DrivingDiffusion [12], and Panacea [36], plays an instrumental role in facilitating the synthesis of images that maintain visual consistency across varying camera perspectives. For additional information, please see the Appendix.
@@
Text Cross-attention enhances Stable Diffusion [24] model’s ability to modulate street scenes through textual scene description. This capability is crucial for dynamically adapting the rendering of street scenes to accommodate various lighting and weather conditions. By integrating detailed textual scene description, PerLDiff can alter visual elements such as illumination and atmospheric effects, ensuring that the generated images reflect the specified conditions. For qualitative examples, see the the Appendix.
3.3 Discussion
In contrast to previous approaches for autonomous driving such as BEVControl [38], MagicDrive [6], DrivingDiffusion [12], and Panacea [36], which employ a basic cross-attention mechanism to integrate controlling condition information, our PerLDiff leverages geometric priors via PerL masking map. This approach directs the generation of each object with its respective control information during the training phase, effectively countering the common misalignment between the attention map and condition information that often results in compromised image controllability. For instance, the attention map of BEVControl [38], illustrated in Fig. 3, demonstrates disorganized patterns and lacks precision. Conversely, our PerLDiff markedly enhances the accuracy of generated images and the granularity of condition information at the object level by ensuring meticulous guidance within the attention map. For more qualitative results, please see the Appendix.
4 Experiments
We assess PerLDiff’s ability to control quality across multiple benchmarks, including multi-view 3D object detection [14, 16, 40, 34], BEV segmentation [41] and monocular 3D object detection [40]. Subsequently, we conduct ablation studies to ascertain the contribution of each component within our proposed methodology.
4.1 Datasets
NuScenes dataset comprises 1,000 urban street scenes, traditionally segmented into 700 for training, 150 for validation, and 150 for testing. Each scene features six high-resolution images (900$\times$1600), which together provide a complete 360-degree panoramic view of the surroundings. Additionally, NuScenes includes comprehensive road maps of the driving environment, featuring details such as lane markings and obstacles. We extend the class and road type annotations similar to MagicDrive [6] and NuScenes, incorporating ten object classes and eight road types for map rendering. To address the resolution limitations of the U-Net architecture in Stable Diffusion [25], we adopt image resolutions of 256$\times$384 as in BEVFormer [14], and 256$\times$704 following BEVFusion (Camera-Only) [16].
KITTI dataset contains 3,712 images for training and 3,769 images for validation. KITTI has only one perspective image and does not have road map information. Given the varied image resolutions in KITTI (approximately 375$\times$1242), we pad them to 384$\times$1280 for generative learning.
4.2 Main Results
In this subsection, we assess PerLDiff’s generative quality through the perception results of several pre-trained methods: BEVFormer [14], BEVFusion (Camera-Only) [16], and StreamPETR [34] for multi-view 3D detection; CVT [41] for BEV segmentation; and MonoFlex [40] for monocular 3D detection trained on the KITTI. Additionally, we leverage synthesized dataset to enhance the performance of various 3D detection models (i.e., BEVFormer and StreamPETR) on the NuScenes test, validating the effectiveness of our PerLDiff.